OpenStudy (anonymous):

Write the first five terms of the sequence defined recursively. Use the pattern to write the nth term of the sequence as a function of n. (Assume that n begins with 1.) a[1]=5,a[k+1]=-a[k]

OpenStudy (anonymous):

5, -5, 5, -5, 5

OpenStudy (anonymous):

SO its a geometric equation I need

OpenStudy (anonymous):

the n th term is (-1)^(n+1) * 5

OpenStudy (anonymous):

the sequence would be f(n) = 5*(-1)^(n+1)

OpenStudy (anonymous):

ok well these are my options: a[n]=5n a[n]=5(-1)^n a[n]=-5^n a[n]=-5^(n-1) a[n]=5(-1)^(n-1)

OpenStudy (anonymous):

the last one. (-1)^(n+1) = (-1)^(n-1)

OpenStudy (anonymous):

oh ok I see now thanks:)

OpenStudy (anonymous):

Find the sum of the infinite series. \[\sum_{i=1}^{\infty}2(-1/4)^i\] A. Undefined B. -2/3 C. 2 D.4/5 E. -2/5

OpenStudy (anonymous):

Ratio test shows that it converges. I cannot recall how to find what it converges to however.

OpenStudy (anonymous):

oh ok thanks...

OpenStudy (anonymous):

This is a geometric series with a = -1/4 and r = -1/4, after you factor the 2 out

OpenStudy (anonymous):

So the answer is E

OpenStudy (anonymous):

The infinite sum for a geometric series is a/(1-r) provided |r| < 1

OpenStudy (anonymous):

oh ok so I would just plug those in and get my answer! Thanks so much!

OpenStudy (anonymous):

you're welcome